# Evaluating LLMs’ Mathematical and Coding Competency through Ontology-guided Interventions

### v4

@@ Abstract
Recent advancements in Large Language Models (LLMs) have showcased striking results on existing logical reasoning benchmarks, with some models even surpassing human performance. However, the true depth of their competencies and robustness in reasoning tasks remains an open question. To this end, in this paper, we focus on two popular reasoning tasks: arithmetic reasoning and code generation. Particularly, we introduce: (i) a general ontology of perturbations for maths and coding questions, (ii) a semi-automatic method to apply these perturbations, and (iii) two datasets, More and Core, respectively, of perturbed maths and coding problems to probe LLM capabilities in numeric reasoning and coding tasks. Through comprehensive evaluations of both closed-source and open-source LLMs, we show a significant performance drop across all the models against the perturbed questions, suggesting that the current LLMs lack robust problem solving skills and structured reasoning abilities in many areas, as defined by our ontology.
@@ 1 Introduction
Logical reasoning in a structured and well-defined domain, such as mathematics and programming, becomes increasingly harder with the increasing presence of interspersed and diverse situations, events, and contexts formulated through natural language queries. Current state-of-the-art Large Language Models (LLM) have shown impressive performance on mathematical problems Cobbe et al. (2021a) and reasonable performance on coding problems Chen et al. (2021a) expressed in natural language. However, these evaluations barely test the depth of LLMs’ expertise, and thus we do not currently have clear insights into the LLM capabilities in these domains. For example, in mathematics, GPT-4’s performance monotonically decreases from GSM-8k Cobbe et al. (2021a) (; 5-shot CoT) on grade school mathematical problems demanding rigorous arithmetic and logical reasoning to solve; to MMLU-Math () (Hendrycks et al., 2020) on a collection of mathematical problems, ranging in difficulty from elementary to advanced levels; and to MATH () (Hendrycks et al., 2021) on challenging competition mathematics problems. Similar variance in LLM performance can also be observed for coding challenges Chen et al. (2021a). Such shallow evaluations are unfit for an objective measure of the finer LLM capabilities as (i) many LLMs like GPT-4 OpenAI (2023) are exposed to publicly available math and coding datasets during pre-training; and ii) many datasets focus on advanced branches of mathematics and problems without bolstering the fundamentals. Hence, before testing the LLMs’ breadth of capabilities by delving into higher mathematics and evaluating competitive coding questions, we instead focus on depth through one fundamental question: {mdframed}[backgroundcolor=green!5] How robust are the capabilities of LLMs in terms of reasoning and understanding of the problem-solving process?
In this work, our goal is to provide an evaluation mechanism that provides clear insights into the robustness of the reasoning abilities of LLMs in the context of maths and coding. Following previous work towards probing language models Ribeiro et al. (2020); Wu et al. (2023); Li et al. (2024a); Wang et al. (2024), we evaluate the robustness of LLMs’ understanding of interesting linguistic and logical structures and derive insights based on them.
@@
Specifically, we design an adaptive dynamic evaluation benchmark through novel ontology-guided perturbations on existing problems. We introduce a novel ontology of perturbation operations that lists various changes across a diverse set of factors, which we apply to previously introduced arithmetic and coding problems. These perturbations allow us to assess whether the model comprehends underlying concepts. For instance, while a model may correctly answer questions in a dataset like GSM8k, it might struggle when presented with a simple perturbation to the question, such as replacing numerical values (Srivastava et al., 2024) in maths questions with variables, which challenges the model to establish relationships among the variables, revealing its deeper understanding (or lack thereof). By introducing these ontological perturbations, (1) we gain insights into the models’ reasoning abilities and (2) uncover strategies for future data augmentation that can then be utilized to enhance LLMs through weakly supervised fine-tuning methodologies.
Our ontology consists of 44 types of perturbations, which we apply to sample questions from GSM8K and coding questions from HumanEval, resulting in and perturbed questions respectively. Our evaluation of GPT-4, GPT-3.5, Metamath, Llama-code, Llama3-Instruct, and Gemini 1.5 shows that most of these models very quickly degrade under different perturbation types. Our contributions are as follows:
@@
We propose a novel, extensive, and extensible ontology of perturbation operations for basic-math- and coding-based reasoning tasks expressed in natural language.
@@
We present a semi-automatic method to exercise such perturbations first through GPT-4, followed by manual filtering. We generate two datasets More and Core—Mathematics- and Code-Oriented Robustness Evaluation, respectively—consisting of maths and coding questions.
@@
We gain insights into the range of capabilities and limitations on such math and coding tasks for several LLMs.
@@ 2 Related Work
MORE-CORE (Our) | math(GSM8K), code(HumanEval) | Dynamic Ontology | GPT-4, Human (Q,A) | 44 |
@@
Therefore, Recent work has focused on the robustness of reasoning under various perturbations that alter reasoning question. Different domain-specific methods have been proposed for generating test cases for reasoning tasks (Yu et al., 2023a; Wu et al., 2023), as summarized in Table 1. In the field of mathematics, contemporary works have employed techniques such as numerical or symbolic substitutions (Li et al., 2024a; Zhou et al., 2023; Meadows et al., 2023; Wang et al., 2024; Patel et al., 2021), the insertion of irrelevant distractors (Shi et al., 2023; Li et al., 2023), functional equivalence (Srivastava et al., 2024), and reverse prediction (Yu et al., 2023b; Berglund et al., 2023; Deb et al., 2023) to uncover conceptual errors (Sanyal et al., 2022), cognitive biases (Dasgupta et al., 2022), or sensitivity to reasoning context (Wu et al., 2023). Similarly, Perturbations are done based on Abstract Syntax Tree guided perturbations are performed on the domain of code (Naik et al., 2023; Allamanis et al., 2021). However, To our knowledge, well-established perturbation methods beyond the domain of mathematics are lacking. In this work, we consolidate and develop a broader underlying ontology that connects and expands upon previous methods for perturbing reasoning datasets. This new framework is both systematic and hierarchical, and it is readily adaptable to various domains, including mathematics and coding.
@@ 3 The Ontology of Perturbations
3.1 The Need for Ontology-based Perturbations
We plan to first identify a set of factors upon which the solution of a structured reasoning problem (expressed in natural language) may depend on (similar to Kaushik et al. (2021)); and perturb a seed question under these set of factors semi-automatically in a model-agnostic way (i.e., not necessarily adversarial to a target model). In the NLI context, Kaushik et al. (2021) utilized human workers to directly perturb a hypothesis, keeping the premise constant; and in a post-hoc way, identifies the categories (or factors) which such revisions pertain to. Previous works Xu et al. (2023); Li et al. (2024a); Wang et al. (2024) discusses ways of perturbation, by identifying a set of factors which is specifically designed to increase the complexity of a seed questions in limited ways. The categories are broad and do not exploit the logical nature of the underlying domain (along with the linguistic dimensions of the instruction). This is where, we believed, an ontological approach may help, where broader categories can help us generalize, while fine-grained sub-categories exploit the domain-specific characteristics.
@@ 3.2 The Ontology
Level IV: Category. This level captures the method (the HOW) through which the higher-level Dimension perturbation is achieved. These methods are domain dependent and, thus, their implementations vary from maths to coding problems.
Aspect (Level I) | Target (Level II) | Dimension (Level III) | Category (Level IV) | Math | Code |
@@
Structural Perturbation Def: Modification on specific aspects of logic or concepts to alter the reasoning process required to reach the answer | Logic Def: Modifications to the reasoning framework or logic underpinning a problem. | Granularity Adjustment Def: Fine-grained sub-tasks of the original question | (iv) Remove Constraint | Remove Constraint | Remove Constraint |
(ii) Partial Solution | Median Inquiry | Helper Function | |||
(vi) Solution Plan | Solution Plan | Solution Plan | |||
(vii) Detail Expansion | Detail Elaboration | Example Detail | |||
@@
Reasoning Adjustment Def: Target at logical structure of the original | (vii) Add Restriction | Restrict Question | Restrict Requirement | ||
G9. Subsequent Question | Further Question | Further Requirement | |||
G10. Concurrent Question | Parallel Question | Parallel Requirement | |||
G11. Change Question | Change Query | Change Docstring | |||
G12. Info Recombination | Info Recombination | Info Recombination | |||
G13. Domain Knowledge | Theoretical Challenge | Code Import | |||
G14. Complex Reality | Value Probability | Example Boundary | |||
G15. General Solution | Code Implementation | Higher Order | |||
@@
Computation Adjustment Def: Target at values or entities | G15. Computation Demand | Value Big | Generalize Parameter | ||
G17. Change Value | Change Subject | Parameter Content | |||
G18. Change Operation | Change Calculation | Variable Type | |||
@@
Formulation Adjustment Def: Reformulate question for solution form to be an abstract expression. | G18. Symbolic Response | Variable Response | Code Execution | ||
G20. Value Relationship | Variable Relation | Parameter Relationship | |||
G21. Variable Group | Variable Scaling | Variable Substitution | |||
G22. Backward Reasoning | Variable Adaptation | Reverse Engineering | |||
G23. Counterfactual | WhatIf Question | WhatIf Code | |||
G24. Solve Value | Solve X | Solve Input | |||
G25. Identify Range | Variable Range | Variable Range | |||
@@
Concept Def: Examination and Analysis of the underlying concepts and principles of a problem | Question Understanding Def: Interpretation of the information inside the question | I.2 Inherent Premise | Identify Assumption | Test Case | |
(i)G24. Complete Missing | Info Sufficiency | Incomplete Answer | |||
(i)G25. Question Formulation | Question Formulation | Question Formulation | |||
(i)G26. Add Misinformation | Introduce Distraction | Introduce Bias | |||
@@
Solution Understanding Def: Assessment of the problem-solving processes | (ii) Optimize Solution | Info Necessity | Reduce Complexity | ||
(ii)G28. Step Functionality | Step Necessity | Step Necessity | |||
(ii)G29. Theoretical Basis | Theoretical Basis | Theoretical Basis | |||
(ii)G30. Cost Analysis | Solution Efficiency | Code Complexity | |||
Critical Thinking Def: Identification of noise, inaccuracies and inconsistencies | (iii) Seek Clarification | Introduce Ambiguity | Example Requirement | ||
(iii)G32. Conditional Analysis | Discuss Separately | Incomplete Requirement | |||
(iii)G33. Conflicting Information | Introduce Contradiction | Wrong Example | |||
(iii)G34. Surface Error | Value Uncommon | Runtime Error | |||
(iii)G35. Hidden Error | Value Error | Logical Error | |||
@@
Representational Perturbation Def: Preservation of the underlying logic and conceptual framework, but modification of the encoding or representation | Question Format Def: Direct modification on the encoding of the question while keeping the logical structure intact | Format Change Def: Rephrasing the question in a different format | I.3 Setting Rephrase | Change Setting | Realworld Usecase |
(i)G37. Change Sequence | Change Sequence | Parameter Sequence | |||
(i)G38. Close Format | True False | True False | |||
(i)G39. Data Restructuring | Value Structuring | Complex Docstring | |||
Format Comparison Def: Comparing two problem of different forms | (i)G40. Identical Problem | Identical Question | Identical Code | ||
Answer Format Def: Indirect modification on the output form | Format Constraint Def: Add constraint on the solution | (ii) Reasoning Format | Binary Coded | No Keyword | |
(ii)G42. Reasoning Style | X Language | X Language | |||
(ii)G43. Alternative Answer | Alternative Answer | Alternative Answer | |||
(ii)G44. New Rule | Define Rules | Simple Name |
4 Curation of More and Core
@@
Our objective is to assess the resilience of LLMs to perturbations of maths and coding questions along various dimensions. Thus, as seed datasets, we use GSM8K (Cobbe et al., 2021b)—a collection of mathematical problems demanding rigorous arithmetic and logical reasoning—and HumanEval Chen et al. (2021a) for coding. Five questions 111Maths questions in the GSM8K dataset take between two and eight steps to solve. We randomly chose five questions that take three to seven steps to solve. We cover various topics involving algebraic questions, physical application questions, and decision-based application questions from GSM8K are perturbed using our ontological framework (see Appendix I) to generate More. On the other hand, we sampled five coding problems from HumanEval dataset (Chen et al., 2021a) that were perturbed using the ontology explained in Appendix I. These perturbations are aimed at modifying the problems in terms of complexity and representation to assess the robustness of the LLMs to these ontological categories of perturbations. Fig. 2 shows examples of three perturbed questions and answers from More and Core. Examples and definitions of all the remaining perturbations are present in Appendix I. We use a three-staged combination of automatic generation from GPT-4 (OpenAI, 2023) with human verification and annotation to create More and Core: (i) perturbed question generation (Section 4.1), (ii) filtering and validation of generated questions (Section 4.2), and (iii) annotating final answers (Section 4.3).

### v5

@@ Abstract
Recent advancements in Large Language Models (LLMs) have showcased striking results on existing logical reasoning benchmarks, with some models even surpassing human performance. However, the true depth of their competencies and robustness in reasoning tasks remains an open question. To this end, in this paper, we focus on two popular reasoning tasks: arithmetic reasoning and code generation. Particularly, we introduce (i) a general ontology of perturbations for math and coding questions, (ii) a semi-automatic method to apply these perturbations, and (iii) two datasets, GSMore and HumanEval-Core, respectively, of perturbed math and coding problems to probe LLM capabilities in numeric reasoning and coding tasks. Through comprehensive evaluations of both closed-source and open-source LLMs, we show a significant performance drop across all the models against the perturbed questions, suggesting that the current LLMs lack robust problem solving skills and structured reasoning abilities in many areas, as defined by our ontology. We open-source the datasets and source codes at: https://github.com/declare-lab/LLM-ReasoningTest.
@@ 1 Introduction
Logical reasoning, that includes mathematics and programming, in a structured and well-defined domain, becomes increasingly complex with the increasing presence of diverse situations, events, and contexts represented in natural language. However, the process of solving a mathematical expression should remain consistent and a logical reasoner should be robust to certain changes in the context. Consider, problems I) I have two apples and Sam has three more. How many apples are there?, and II) I see two wombats. And my friend sees another three. How many wombats are there?. Variations I and II should yield the same expression (), and a similar process to solve resulting in the same answer (). Current state-of-the-art Large Language Models (LLM) have shown impressive performance on a wide variety of mathematical problems Cobbe et al. (2021a) and coding problems Chen et al. (2021a) expressed in natural language. However, these evaluations barely test the depth of LLMs’ expertise, and thus we do not currently have fine insights into the LLM capabilities in these domains. For example, in mathematics, GPT-4’s performance monotonically decreases from GSM-8k Cobbe et al. (2021a) (; 5-shot CoT) on grade school mathematical problems requiring rigorous arithmetic and logical reasoning to solve; to MMLU-Math () (Hendrycks et al., 2020) on a collection of mathematical problems, ranging in difficulty from elementary to advanced levels; and to MATH () (Hendrycks et al., 2021) on challenging competitive mathematics problems. Similar variance in LLM performance can also be observed for coding challenges Chen et al. (2021a). Such high-level evaluations show the breadth of LLM capabilities, but do not provide any insights into the finer and more fundamental math- and coding-relevant capabilities and shortcomings of LLMs as (i) many LLMs like GPT-4 OpenAI (2023) are exposed to publicly available math and coding datasets during pre-training, and (ii) many datasets focus on advanced branches of mathematics and problems, skipping the fundamentals. Hence, we instead focus on depth, encapsulated by one fundamental question:
@@
Evaluating LLMs on these modified questions could provide fine-grained insights into the robustness of the reasoning abilities in the context of math and coding. Specifically, our evaluation benchmark is based on our novel ontology of perturbation operations that lists various changes across a diverse set of factors. We apply these perturbations to existing arithmetic and coding problems. These perturbations allow us to assess whether the model truly comprehends underlying concepts and structures or regurgitates. For instance, while a model may correctly answer a math question in a dataset like GSM8k, it may struggle with a simple perturbation to the question, such as replacing numerical values (Srivastava et al., 2024) in math questions with variables, which challenges the model to establish relationships among the variables, revealing its deeper understanding (or lack thereof). By introducing these ontological perturbations, (1) we gain insights into the models’ reasoning abilities and (2) uncover strategies for future data augmentation that can then be utilized to enhance LLMs through weakly supervised fine-tuning methodologies.
Our perturbations extend beyond mere numerical alterations to seed questions; they also challenge with the underlying concepts and principles required to solve a problem. For instance, we introduce a category Critical Thinking, which includes scenarios where multiple correct answers may be possible, depending on the constraints on various variables, as demonstrated in the example below:
Such instances are unique to our datasets making the evaluation more rigorous.
We sample five questions from GSM8K, subjecting them to 44 types of perturbations, followed by extensive manual filtering and validation, resulting in perturbed questions in GSMore. As our ontology is generic, we also apply these perturbations to randomly sampled five code-generation questions from HumanEval producing questions in HumanEval-Core. As a whole, this work makes the following contributions:
@@
We propose a novel, extensive, and extensible ontology of perturbation operations for math and coding problems.
@@
We provide a way to semi-automatically apply such perturbations first through GPT-4, followed by manual filtering. We refer to this dataset as GSMore— Mathematics-oriented Robustness Evaluation, consisting of examples starting from just five questions in GSM8K.
@@
We show that the ontology and the semi-automatic pipeline can be easily extended to domains like code-generation, by creating the dataset HumanEval-Core — Coding-oriented Robustness Evaluation, having samples from seed questions in HumanEval.
4.
We further evaluate the math and code-based reasoning abilities of the latest LLMs, including GPT-4o and o1, using our datasets, revealing their vulnerability to these perturbations.
5.
The integration of our ontology with the newly introduced dataset, referred to as GSMore and HumanEval-Core, paves the way for a fresh perspective in assessing the math and code reasoning capabilities of LLMs. This evaluation approach transcends the often unidimensional leaderboard performance metrics by delving into LLM performance on finely defined classes of problems.
@@ 2 Related Work
| GSMore-HumanEval-Core (Our) | math (GSM8K), code (HumanEval) | Dynamic Ontology | GPT-4, Human (Q,A) | 44 |
@@
Therefore, Recent work has focused on the robustness of reasoning under various perturbations that alter reasoning questions. Different domain-specific methods have been proposed for generating test cases for reasoning tasks (Yu et al., 2023a; Wu et al., 2023), as summarized in Table 1. In the field of mathematics, contemporary works have employed techniques such as numerical or symbolic substitutions (Li et al., 2024a; Zhou et al., 2023; Meadows et al., 2023; Wang et al., 2024; Patel et al., 2021), the insertion of irrelevant distractors (Shi et al., 2023; Li et al., 2023), functional equivalence (Srivastava et al., 2024), and reverse prediction (Yu et al., 2023b; Berglund et al., 2023; Deb et al., 2023) to uncover conceptual errors (Sanyal et al., 2022), cognitive biases (Dasgupta et al., 2022), or sensitivity to reasoning context (Wu et al., 2023). Similarly, Abstract Syntax Tree guided perturbations are performed on the domain of code (Naik et al., 2023; Allamanis et al., 2021). However, to our knowledge, perturbation methods across the domain of reasoning tasks are lacking. In this work, we consolidate and develop a broader underlying ontology that connects and expands upon previous methods for perturbing reasoning datasets. This new framework is both systematic and hierarchical, and it is readily adaptable to various domains, including mathematics and coding.
@@ 3 The Ontology of Perturbations
3.1 Need for Ontology of Perturbations
We first identify a set of factors relevant to the solution of a structured reasoning problem expressed in natural language (similar to Kaushik et al. (2021)); and perturb a seed question under these set of factors semi-automatically in a model-agnostic way (i.e., not necessarily adversarial to a target model). In the NLI context, Kaushik et al. (2021) utilized human workers to directly perturb a hypothesis, keeping the premise constant; and in a post-hoc way, identifies the categories (or factors) which such revisions pertain to. Previous works Xu et al. (2023); Li et al. (2024a); Wang et al. (2024) discusses ways of perturbation, by identifying a set of factors which is specifically designed to increase the complexity of a seed questions in limited ways. The categories are broad and do not exploit the logical nature of the underlying domain (along with the linguistic dimensions of the instruction). This is where, we believed, an ontological approach may help, where broader categories can help us generalize, while fine-grained sub-categories exploit the domain-specific characteristics.
@@ 3.2 The Ontology
Level IV: Category. This level captures the method (the HOW) through which the higher-level Dimension perturbation is achieved. These methods are domain dependent and, thus, their implementations vary from math to coding problems.
l|l|l|l|l|l
Aspect (Level I) Target (Level II) Dimension (Level III) Category (Level IV) Math Code
Structural Perturbation
@@
Def: Modification on specific aspects of logic or concepts to alter the reasoning process required to reach the answer Logic Alteration
Def: Modifications to the reasoning framework or logic underpinning a problem.
Granularity Adjustment
Def: Fine-grained sub-tasks of the original question
(iv) Remove Constraint Remove Constraint Remove Constraint
(v) Partial Solution Median Inquiry Helper Function
(vi) Solution Plan Solution Plan Solution Plan
(vii) Detail Expansion Detail Elaboration Example Detail
Reasoning Adjustment
@@
Def: Target at logical structure of the original G8. Add Restriction Restrict Question Restrict Requirement
G9. Subsequent Question Further Question Further Requirement
G10. Concurrent Question Parallel Question Parallel Requirement
G11. Change Question Change Query Change Docstring
G12. Info Recombination Info Recombination Info Recombination
G13. Domain Knowledge Theoretical Challenge Code Import
G14. Complex Reality Value Probability Example Boundary
G15. General Solution Code Implementation Higher Order
Computation Adjustment
@@
Def: Target at values or entities G16. Computation Demand Value Big Generalize Parameter
G17. Change Value Change Subject Parameter Content
G18. Change Operation Change Calculation Variable Type
Formulation Adjustment
@@
Def: Reformulate question for solution form to be an abstract expression. G19. Symbolic Response Variable Response Code Execution
G20. Value Relationship Variable Relation Parameter Relationship
G21. Variable Group Variable Scaling Variable Substitution
G22. Backward Reasoning Variable Adaptation Reverse Engineering
G23. Counterfactual WhatIf Question WhatIf Code
G24. Solve Value Solve X Solve Input
G25. Identify Range Variable Range Variable Range
Concept Analysis
Def: Examination and Analysis of the underlying concepts and principles of a problem
Question Understanding
@@
Def: Interpretation of the information inside the question (i)G23. Inherent Premise Identify Assumption Test Case
(i)G24. Complete Missing Info Sufficiency Incomplete Answer
(i)G25. Question Formulation Question Formulation Question Formulation
(i)G26. Add Misinformation Introduce Distraction Introduce Bias
Solution Understanding
@@
Def: Assessment of the problem-solving processes (ii)G27. Optimize Solution Info Necessity Reduce Complexity
(ii)G28. Step Functionality Step Necessity Step Necessity
(ii)G29. Theoretical Basis Theoretical Basis Theoretical Basis
(ii)G30. Cost Analysis Solution Efficiency Code Complexity
Critical Thinking
Def: Identification of noise, inaccuracies and inconsistencies
(iii)G31. Seek Clarification Introduce Ambiguity Example Requirement
(iii)G32. Conditional Analysis Discuss Separately Incomplete Requirement
(iii)G33. Conflicting Information Introduce Contradiction Wrong Example
(iii)G34. Surface Error Value Uncommon Runtime Error
(iii)G35. Hidden Error Value Error Logical Error
Representational Perturbation
Def: Preservation of the underlying logic and conceptual framework, but modification of the encoding or representation
Question Format
@@
Def: Direct modification on the encoding of the question while keeping the logical structure intact Format Change
Def: Rephrasing the question in a different format (i)G36. Setting Rephrase Change Setting Realworld Usecase
(i)G37. Change Sequence Change Sequence Parameter Sequence
(i)G38. Close Format True False True False
(i)G39. Data Restructuring Value Structuring Complex Docstring
Format Comparison
Def: Comparing two problem of different forms (ii)G40. Identical Problem Identical Question Identical Code
Answer Format
Def: Indirect modification on the output form
Format Constraint
Def: Add constraint on the solution (i)G41. Reasoning Format Binary Coded No Keyword
(i)G42. Reasoning Style X Language X Language
(i)G43. Alternative Answer Alternative Answer Alternative Answer
(i)G44. New Rule Define Rules Simple Name
4 Curation of GSMore and HumanEval-Core
@@
Our objective is to assess the resilience of LLMs to perturbations of math and coding questions along various dimensions. Thus, as seed datasets, we use GSM8K (Cobbe et al., 2021b)—a collection of mathematical problems demanding rigorous arithmetic and logical reasoning—and HumanEval Chen et al. (2021a) for coding. Five questions 111Math questions in the GSM8K dataset take between two and eight steps to solve. We randomly chose five questions that take three to seven steps to solve. We cover various topics involving algebraic questions, physical application questions, and decision-based application questions. from GSM8K are perturbed using our ontological framework (see Appendix I) to generate GSMore. On the other hand, we sampled five coding problems from HumanEval dataset (Chen et al., 2021a) that were perturbed using the ontology explained in Appendix I. These perturbations are aimed at modifying the problems in terms of complexity and representation to assess the robustness of the LLMs to these ontological categories of perturbations. Fig. 2 shows examples of three perturbed questions and answers from GSMore and HumanEval-Core. Examples and definitions of all the remaining perturbations are present in Appendix I. We use a three-staged combination of automatic generation from GPT-4 (OpenAI, 2023) with human verification and annotation to create GSMore and HumanEval-Core: (i) perturbed question generation (Section 4.1), (ii) filtering and validation of generated questions (Section 4.2), and (iii) annotating final answers (Section 4.3).